OpenStudy (anonymous):

having trouble figuring out how to solve this. e^(x-5)=(1/e^4)^(x+2)

OpenStudy (anonymous):

Try taking the ln of each side so you can get the exponents down.

OpenStudy (anonymous):

so we have (x-5)ln e=(x+2)ln (1/4e^4) after this what is our next step?

OpenStudy (turingtest):

you have to be careful in simplification here\[(x^a)^b=x^{ab}\]and\[\frac1{x^a}=x^{-a}\]so\[\large (\frac1{e^4})^{x+2}=(e^{-4})^{x+2}=e^{-4(x+2)}\]now try with the log...

OpenStudy (anonymous):

just got it. thx. mental block

OpenStudy (anonymous):

-3/5?